Hiram’s Lakeside is a popular restaurant located on Lake Washington in Seattle. The owner of the restaurant has been trying to better understand costs at the restaurant and has hired a student intern to conduct an activity-based costing study. The intern, in consultation with the owner, identified three major activities and then completed the first-stage allocations of costs to the activity cost pools. The results appear below. Activity Cost Pool Activity Measure Total Cost Total Activity Serving a party of diners Number of parties served $ 11,960 5,200 parties Serving a diner Number of diners served $ 107,010 12,300 diners Serving drinks Number of drinks ordered $ 26,750 10,700 drinks The above costs include all of the costs of the restaurant except for organization-sustaining costs such as rent, property taxes, and top-management salaries. Some costs, such as the cost of cleaning the linens that cover the restaurant's tables, vary with the number of parties served. Other costs, such as washing plates and glasses, depend on the number of diners served or the number of drinks served. Prior to the activity-based costing study, the owner knew very little about the costs of the restaurant. She knew that the total cost for the month (including organization-sustaining costs) was $180,000 and that 12,000 diners had been served. Therefore, the average cost per diner was $15. Required: 1&2. According to the activity-based costing system, what is the total cost and average cost per diner for serving each of the following parties of diners? (Round your intermediate calculations to 2 decimal places. Round your Total Cost final answers to 2 decimal places and your Average Cost final answers to 3 decimal places.)

Total Cost Average Cost
a. A party of four diners who order five drinks in total. per diner
b. A party of three diners who do not order any drinks. per diner
c. A party of one diner who orders four drinks. per diner
Answer #1

Activity rate = Activity cost/total activity

Serving a party = 11960/5200 =$2.3 per party

Serving a diner = 107010/12300 =$8.7 per diner

Serving drinks = 26750/10700 =$2.5 per drink

Hence, cost will be as follows

A. = 2.3 +8.7*4+2.5*5 =$49.6

Average cost per diner =49.6/4 =$12.4

B

Total cost = 2.3 +8.7*3 +0 =$28.4

Average cost per diner =28.4/3 =$9.467

C.

Total cost = 2.3+8.7+2.5*4 =$21

Average cost =$21